[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Sun Apr 1 12:50:03 EDT 2018

Logentry Text:
Yesterday around 3 pm we completed the data taking at 8 deg with SHMS, this being the last kinematic point in the EMC effect experiment run plan. With that we actually completed the data taking for the E12-10-002/E12-10-008 experiments. For a summary of the data we took please check:

SHMS --> https://docs.google.com/document/d/13i-SQGGES52d8fjJo9CTdgMc0A7J7z_t9FdKewW6HEM/edit?ts=5a8db865# 
HMS --> https://docs.google.com/document/d/1scxBFrrasplcp8-AznKRal5rr_19KKGk7Wfwsld-8Fw/edit#heading=h.8xbkknkbe17o

Before transitioning to E12-06-107 (CT) we took some optics data with HMS to define a field correction for the HMS dipole saturation at 5.539 GeV/c (this is the HMS momentum setting for CT). 

While we were taking optics data with HMS we changed SHMS magnets settings (polarity and momentum) and angle to prepare for the CT run. We also switched the DAQ from singles to coincidence mode (in between optics runs when beam was down due to the LCW issue). 

The optics data taking was completed around 10:30 pm. We then started production for CT on C 6%. During owl we completed the C 6% running (both coincidence and singles) and we are now taking data on C 1.5 %. We got about 25 % of the C 1.5 % statistics in about 4 hours (beam was fairly good). The expectation is to run for 12 more hours on this target. Then we would do short Al dummy and LH2 runs. We expect to complete the CT data taking by 8:30 am tomorrow (run plan is posted here: 
https://hallcweb.jlab.org/wiki/index.php/Main_Page#The_10.6_GeV_Run_Plan_.28282_hrs.29 )

Tomorrow beam studies will start at 8:30 am. I was told by the PD this morning that one of the items on their list is to attempt to address the issue with high current delivery to Hall C.

There is some need for people to staff Hall C shifts starting with next weekend. However, we are not so desperate to resort to limericks to inspire people to sign up :) (I, for instance, use far coarser methods).  If you're wondering what's this about check out today's Hall *A* RC daily update :). So to all: take a look at the Hall A (yes! Hall A!) Shift Schedule (https://misportal.jlab.org/mis/physics/shiftSchedule/?experimentRunId=TRITON-2017) and after you recover from your shock *please* do not wait another minute to sign up for Hall *A* shifts.  

There will be no meeting today. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3553010
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180401/79852409/attachment.html>

More information about the Hallc_running mailing list